A plane flew 600 kilometers along a line 35° east of south. How far east of its starting point is it? How far south? Represent the given vectors.


1
Expert's answer
2021-02-24T12:50:02-0500

We can find distance that the plane travels east from its starting point from the geometry:


"cos\theta=\dfrac{d_x}{d},""d_x=dcos\theta=600\ km\cdot cos35^{\circ}=491.5\ km."

We can find distance that the plane travels south from its starting point from the geometry:


"sin\theta=\dfrac{d_y}{d},""d_y=dsin\theta=600\ km\cdot sin35^{\circ}=344.14\ km."